OpenStudy (anonymous):

Medal Awarded-The equation of line EF is y = 1 over 2x + 6. Write an equation of a line parallel to line EF in slope-intercept form that contains point (0, −2).

OpenStudy (anonymous):

y = −2x − 2 y = negative 1 over 2x + 2 y = 1 over 2x − 2 y = − 2x + 2

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@Nicksl99

OpenStudy (anonymous):

@ninjasandtigers

OpenStudy (anonymous):

@Elsa213

OpenStudy (anonymous):

@uri

OpenStudy (anonymous):

@kohai

OpenStudy (anonymous):

do you know how to do this?

OpenStudy (anonymous):

i know your supposed to use y=mx+b

OpenStudy (anonymous):

@Jesstho.-.

OpenStudy (anonymous):

@Reilly208

OpenStudy (anonymous):

@vksmith7

OpenStudy (anonymous):

@All_Hail_Moses

OpenStudy (anonymous):

this is my worst subject

OpenStudy (anonymous):

i need all the help i can get

OpenStudy (anonymous):

hey moses

OpenStudy (anonymous):

are you good at this stuff

OpenStudy (anonymous):

I have no clue bro. Sorry I cheated in Algebra and Geometry.

OpenStudy (anonymous):

k np

OpenStudy (anonymous):

elsa can you help please

Elsa213 (elsa213):

i think the answer is y = − 2x + 2

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

its going to have the same slope as the first equation, so since the slope to the first equation is 1/2x, the second slope will be 1/2x @Freeflex123